OpenStudy (anonymous):

The measure of an angle is 24degrees less than its complement, then the angle is ____.

hartnn (hartnn):

complimentary angles add upto 90 degrees

OpenStudy (anonymous):

yes, please help me

hartnn (hartnn):

so if the measure of the angle is x, what is the measure of its compliment

OpenStudy (anonymous):

it must be 90-x? right?

hartnn (hartnn):

yes!

hartnn (hartnn):

now x is 24 degrees less than 90-x can u set up an equation using this ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

90-24=66?

hartnn (hartnn):

huh? nopes hint : 'now x is 24 degrees less than 90-x' means if u add 24 to x, u get 90-x

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

you mean x+24=90-x

OpenStudy (anonymous):

but it is wrong

hartnn (hartnn):

thats correct, why u say its wrong ?

OpenStudy (anonymous):

2x=90-24=66 x=33? right?

hartnn (hartnn):

yes, absolutely correct!

OpenStudy (anonymous):

thanx a lot champion

hartnn (hartnn):

welcome ^_^